40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2708     Entry time: Wed Mar 24 12:38:17 2010
Author: Hartmut 
Type: Configuration 
Category: Green Locking 
Subject: Broadband PD for green PLL 

Modified one of the PD assemblies carrying a large SI-Diode (~10mm diameter).

Removed elements used for resonant operation and changed PD readout to transimpedance

configuration. The opamp is a CLC409 with 240 Ohm feedback (i.e. transimpedance) resistor.

To prevent noise peaking at very high frequencies and get some decoupling of the PD,

I added a small series resistor in line with the PD and the inverting opamp input.

It was chosen as 13 Ohm, and still allows for operation up to ~100MHz.

Perhaps it could be smaller, but much more bandwith seems not possible with this opamp anyway.

Changes are marked in the schematic, and I list affected components here.

(Numbers refer to version 'PD327.SCH' from 30-April-1997):

-removed L4

-connected L3 (now open pad) via 100 Ohm to RF opamp output. This restores the DC sognal output.

-removed c17

-connected pin 3 of opamp via 25 Ohm to GND

-connected kathode of PD via 13 Ohm to pin 2 of opamp

-removed L6, C26, L5, C18, and C27

-shorted C27 pad to get signal to the RF output


Measured the optical TF with the test laser setup.

(Note that this is at 1064nm, although the PD is meant to work with green light at 532nm!)

Essentially it looks usable out to 100MHz, where the gain dropped only by about

6dB compared to 10MHz.

Beyond 100MHz the TF falls pretty steeply then, probably dominated by the opamp.


The maximal bias used is -150V.

If the bias is 'reduced' from -150V to -50V, the response goes down by 4dB at 10MHz and

by 9dB at 100MHz.

 The average output was 30mV at the RF output, corresponding to 60mV at the opamp output (50Ohm divider chain).

With 240 Ohm transimpedance this yields 250µA photo-current used for these transfer functions.






ELOG V3.1.3-